OpenStudy (anonymous):

Help!! Fan and Medal :) I have to solve the inequality 2/3(x - 12)

OpenStudy (anonymous):

This is what I have so far 2/3x - 8 <or equal to x + 8

OpenStudy (anonymous):

@cj49 @mathmath333

OpenStudy (anonymous):

8... 4+4+4=12 so 2/3 of 12 = 4+4 =8

OpenStudy (anonymous):

im pretty sure thats not how you solve it...thats not how I've solved any of the other problems...

OpenStudy (anonymous):

\[2/3(x - 12) \le x + 8\]

OpenStudy (anonymous):

thats the problem....

OpenStudy (cj49):

@Jumperman

OpenStudy (jumperman):

Well, Inequalities are hard. Let's see if I can remember how to do them.

OpenStudy (anonymous):

2/3(x−12)≤x+8 2/3x - 8 ≤ x + 8 Thats as far as I've gotten...I just distributed 2/3

OpenStudy (anonymous):

idk y it came out like that its supposed to be less than or equal to sign

OpenStudy (mathmath333):

\[0 \le 3x^2+16x-66 \] is coming

OpenStudy (jumperman):

Okay, good.

OpenStudy (anonymous):

thats the next step or your saying thats the answer ? @mathmath333

OpenStudy (jumperman):

Is this a multistep inequality?

OpenStudy (anonymous):

yes @Jumperman

OpenStudy (jumperman):

https://www.youtube.com/watch?v=y7QLay8wrW8 Great video for this.

OpenStudy (cj49):

multiply by 3 2x-8 <or equal to 3x + 24 add 8 2x<or equal to 3x +32 m stucked after this

OpenStudy (mathmath333):

this cant be factored

OpenStudy (anonymous):

ummm ok....im confused :/

OpenStudy (jumperman):

Watch the video, it might help.

OpenStudy (anonymous):

ok I will @Jumperman

OpenStudy (mathmath333):

x>=2.66666666

OpenStudy (anonymous):

how did u get that?.... and should I write it as a fraction?

OpenStudy (jumperman):

2/3 equals 2.66666666.

OpenStudy (jumperman):

I'm wrong sorry.

OpenStudy (mathmath333):

generally they give equations that factors

OpenStudy (mathmath333):

i mean inequalities

OpenStudy (anonymous):

ok.... I put in the equation and they said the answer was x>or equal to -48

OpenStudy (anonymous):

but I dont get how they got the answer

OpenStudy (mathmath333):

either the question or the answer is wrong

OpenStudy (anonymous):

I'm not sure...I'll just ask my teacher....thanks for trying! :) @mathmath333 @cj49 @Jumperman